OpenStudy (anonymous):

PART 1: solve the inequality 3(x+6) >=4x-2 and show all of your work. You can use >= to represent the greater than or equal tp symbol. PART 2: Use complte sentences to describe the graph of the solution from part one. 2nd problem: PART 1: Use complete sentences to decribe a real-world scenario that could be represented by the inequality 5x+2y>or equal to 45. PART 2: Choose one ordered pair that is a solution th the givin inequality and explain what that orderd pair means in the context of your real-world scenario... I really dont get any of these, what so ever.. Help would be great!!!

OpenStudy (anonymous):

help!!!!!!!!

OpenStudy (anonymous):

Part 1; \[3x+18 \ge 4x-2\] add (-3x in both sides) \[3x+18-3x \ge 4x-2-3x\] \[18 \ge x-2\] add(2) in both sides \[18+2 \ge x\] \[20 \ge x\] And that's equal to \[x \le 20\] I guess yo can write sentences on how to graph it

OpenStudy (anonymous):

u da bomb ty